# Image Complexity-Aware Adaptive Retrieval for Efficient Vision-Language Models

### v1

@@ 4.2 Experimental Setup
We evaluate four ICAR variants (ICAR-8/24, ICAR-12/24, ICAR-16/24, ICAR-20/24) indicating early exit and total layers, using ICC complexity-based routing where simple images use early exits while complex images process through all layers. We train on MS-COCO [lin2014microsoft] (82K) and Flickr30K [young2014flickr30k] (29K), then evaluate on both datasets augmented with 100K LAION samples to test web-scale performance. As a web-scraped dataset, LAION provides the rich complexity distribution needed to evaluate adaptive computation under realistic deployment conditions.
ICAR models fine-tune from OpenCLIP ViT-L/14 using AdamW [loshchilov2019decoupled] with differential learning rates (1e-6 backbone, 1e-4 projection heads) for 10 epochs. We measure both retrieval quality and computational efficiency: instance-level retrieval uses Recall@k (exact image-caption matching), category-level retrieval uses mAP@k (broader semantic matching), and efficiency uses practical throughput on A100 GPUs.
@@ 5.1 Retrieval Performance and Key Insights
| Instance Retrieval - Text to Image - R@k | |||||||
@@
| ViT-L/14 (FT) | 71.0 | 89.9 | 93.8 | 48.5 | 73.5 | 81.8 | 175.33 |
| Category Retrieval - Text to Image - mAP@k | |||||||
@@
| ViT-L/14 (FT) | 84.3 | 70.2 | 64.1 | 92.6 | 86.2 | 82.2 | 175.33 |
Instance vs category: Instance and category retrieval have different computational requirements. Instance-level retrieval shows expected performance decline with early exits; for ICAR-8/24, the aggregate recall scores across Flickr+LAION and COCO+LAION is 95.0% of the baseline111Performance retention calculated using aggregate RSUM [li2022image] methodology: sum of all ICAR-8/24 retrieval scores divided by sum of baseline scores across both datasets.. For category-level retrieval, ICAR variants generally maintain or exceed baseline performance with small, dataset-specific differences: on COCO+LAION, mAP@10 can be marginally higher (e.g., 93.1 vs 92.6 baseline), whereas on Flickr+LAION, ICAR variants remain within 1–2 points of baseline across mAP@k.
Hypothesis validation: The results support our hypothesis about the progressive nature of transformer processing: early layers capture coarse semantic features adequate for category matching, while deeper layers refine instance-level discriminative details. Our dual-path training acts as implicit regularisation, creating more robust categorical representations through multi-scale supervision.
Regularisation effect: The maintained category-level performance is consistent with dual-path training acting as an implicit regularisation technique. Multi-scale supervision from both early and full path losses helps learn robust categorical representations. Early layers focus on coarse semantic features sufficient for category matching, while deeper layers refine instance-level details.
@@ 5.2 Efficiency Analysis and Impact
ICAR achieves 20% practical speedup (ICAR-8/24: 1.195, 120.0 vs 100.4 images/second) with GFLOP-estimated potential up to 1.41. The gap between estimated and measured performance requires production-grade optimisations to fully realise GFLOP reduction potential. Figure 6(a) compares estimated versus measured A100 speedups, while Figure 6(b) shows differential impact: category-level retrieval maintains or exceeds baseline (ICAR-8/24: 100.1% retention), whereas instance-level shows expected degradation. This supports ICAR’s strategy: category matching requires only coarse semantic features from intermediate layers, while instance matching demands fine-grained features from deeper processing.
Production-Scale Impact: If efficiency improvements such as ICAR’s 20% were implemented at big tech production scale, processing Google Photos’ 6 billion daily images [GoogleIO2024] could save 667 A100 GPU-hours daily and 133,640 kWh annually, equivalent to eliminating 16.6 metric tonnes of CO2 emissions222Projected calculation: Google Photos processes 6 billion images daily [GoogleIO2024]. Production-grade optimisations achieve 3-7× speedups [NVIDIADocsOptimization], suggesting 500 images/second on A100 GPUs. Daily processing requires 3,333 GPU-hours (6 billion ÷ 500 ÷ 3,600 seconds/hour). With A100’s 500W consumption and datacenter PUE of 1.10 [GoogleDataCenters], annual consumption is 668,002 kWh. ICAR’s 20% reduction could save 667 GPU-hours daily and 133,640 kWh annually. CO2 calculation uses UK grid intensity [CarbonBrief2024]..
@@ 6 Conclusion
This paper presented ICAR (Image Complexity-Aware Retrieval), the first adaptive computation method for image-text retrieval that addresses the cross-modal alignment challenge through unified embedding spaces. At the core of ICAR is ConvNeXt-IC, a lightweight image complexity classifier that serves as the routing mechanism: it predicts whether an image is simple or complex and routes simple images to early exits while complex images use full processing. During training, both the early-exit and full paths are jointly supervised via dual-path contrastive learning so that all vision embeddings remain compatible with the text embedding in a single unified space. By recognising complexity assessment as a classification task, ConvNeXt-IC achieves state-of-the-art performance (0.959 Pearson correlation, 4.4 speedup), and ICAR’s dual-path training enables 20% practical speedup with maintained retrieval quality.
Future work could enhance the approach through dynamic exit selection based on intermediate confidence scores, end-to-end integration of complexity assessment within the vision-language learning objective, and production-grade optimisations to bridge the gap between GFLOP-estimated (1.41×) and measured (1.195×) speedup results.
6.0.1 Acknowledgements
To facilitate reproducibility and future research, all code, trained model weights, and the annotated LAION complexity dataset are publicly available at https://github.com/MikelWL/ICAR.

### v2

@@ 4.2 Experimental Setup
Datasets: MS-COCO [20] and Flickr30K [42] are standard image–text retrieval benchmarks with curated image-caption pairs. To evaluate retrieval under conditions that are more representative of real-world data, the retrieval pool is augmented using LAION-COCO [32]. LAION [32] is a large-scale dataset of image–text pairs that was collected through automation from publicly available web sources, and LAION-COCO is a subset of it which contains text in the same caption style as MS-COCO and Flickr30K. Training uses the MS-COCO (82K) and Flickr30K (29K) training sets, and evaluation uses the respective test sets (5K test set for MS-COCO and 1K test set for Flickr30K) each augmented with an additional 100K LAION-COCO samples.
Variants and baseline: ICAR-x/24 denotes an early exit after layer x, out of 24 vision layers. Four variants are evaluated with x=8, 12, 16, and 20 (ICAR-8/24, ICAR-12/24, ICAR-16/24, ICAR-20/24), covering compute–quality trade-offs from early to late exits. ViT-L/14 (OpenCLIP) is used as the baseline because ICAR is built directly on top of the same backbone [16]. This isolates the effect of ICC-based routing and early exits.
Training: ICAR is fine-tuned from OpenCLIP ViT-L/14 [16] using AdamW [22] with differential learning rates (1e-6 backbone, 1e-4 projection heads) for 10 epochs. Weight decay is set to 0.01 with gradient clipping (1.0), cosine annealing with warmup, and equal weighting of early and full losses (). Training uses a batch size of 128, and the best validation checkpoint across epochs is used for evaluation.
Tasks and metrics: We evaluate text-to-image retrieval under two settings. For instance-level retrieval, each query has a single correct image, and retrieval quality is measured using Recall@k (R@k). For category-level retrieval, we map samples to MS-COCO semantic categories and treat retrievals as correct when the query and retrieved image belong to the same category, and retrieval quality is measured using mean Average Precision@k (mAP@k). GFLOPs are reported as a weighted average over the routing distribution (49.5% simple, 50.5% complex), accounting for the chosen early-exit depth and ICC overhead. Lower GFLOPs indicate less compute per sample and are therefore preferable for efficiency.
@@ 5.1 Retrieval Performance and Key Insights
Tables 3 and 4 show the trade-off between early exiting and retrieval quality across instance-level and category-level retrieval.
| Instance Retrieval (R@k) | |||||||
@@
| [16] | 71.0 | 89.9 | 93.8 | 48.5 | 73.5 | 81.8 | 175.33 |
| Category Retrieval (mAP@k) | |||||||
@@
| [16] | 84.3 | 70.2 | 64.1 | 92.6 | 86.2 | 82.2 | 175.33 |
Instance-level retrieval: Instance-level retrieval evaluates exact matching between specific image–text pairs, which places more emphasis on fine-grained details. In our setup, the additional LAION-COCO samples primarily act as distractor candidates for these queries, so routing simple candidates through earlier exits mainly affects efficiency, and only harms quality if early-exit embeddings become insufficiently discriminative. Overall, instance-level retrieval shows a gradual performance decline as the exit layer becomes earlier; for ICAR-8/24, the aggregate recall scores across Flickr+LAION and COCO+LAION are 95.0% of the baseline111Performance retention calculated using aggregate RSUM [18] methodology: sum of all ICAR-8/24 retrieval scores divided by sum of baseline scores across both datasets..
Category-level retrieval: Category-level retrieval treats any candidate from the same category as relevant, so simple web images can become valid matches for broader queries. This makes category-level performance a direct test of whether early-exit embeddings preserve coarse semantic information well enough for matching. In Table 4, ICAR variants generally maintain category-level performance with small, dataset-specific differences: on COCO+LAION, mAP@10 can be marginally higher (e.g., 93.1 vs 92.6 baseline), whereas on Flickr+LAION, ICAR variants remain within 1–2 points of baseline across mAP@k.
Coarse-to-fine features: The results support our hypothesis about the progressive nature of transformer processing: early layers capture coarse semantic features that are often sufficient for category matching, while deeper layers refine instance-level discriminative details. The strong category-level performance is consistent with dual-path training providing multi-scale supervision that acts as implicit regularisation, improving robustness of early-exit embeddings for coarse matching.
@@ 5.2 Efficiency Analysis and Impact
Routing coverage: Most early exits come from the LAION-COCO portion of the retrieval pool rather than the benchmark test images. ICC routes 195/5,000 MS-COCO test images (3.9%) and 2/1,000 Flickr30K test images (0.2%) to early exits, but routes 49,314/100,000 LAION-COCO candidates (49.3%). This indicates that most compute savings come from routing web candidates, while curated benchmark images are rarely routed to early exits. This behaviour is desirable for our target use case: curated benchmarks tend to contain more complex images, whereas web data includes a broader mix of simple and complex samples. ICAR therefore targets the part of the retrieval pool where there is the most opportunity for compute savings.
Measured vs estimated speedups: ICAR achieves 20% faster image encoding (ICAR-8/24: 1.195, 120.0 vs 100.4 images/second) with GFLOP-estimated potential up to 1.41. The gap between estimated and measured speedups reflects additional overheads beyond raw transformer compute, including routing, framework and kernel overheads, so GFLOP reduction is an upper bound on throughput gains. Figure 6(a) compares estimated versus measured A100 speedups, while Figure 6(b) shows differential impact: category-level retrieval maintains or exceeds baseline (ICAR-8/24: 100.1% retention), whereas instance-level shows expected degradation.
Production-scale impact: If efficiency improvements such as ICAR’s 20% were implemented at big tech production scale, processing Google Photos’ 6 billion daily images [12] could save 667 A100 GPU-hours daily and 133,640 kWh annually, equivalent to eliminating 16.6 metric tonnes of CO2 emissions222Projection (A100): Google Photos processes images/day [12]. Assume 500 img/s per GPU [24]. GPU-hours/day . Annual energy kWh (PUE 1.10) [11]. A 20% reduction saves 667 GPU-hours/day and 133,640 kWh/year. CO2 uses UK grid intensity [3] ( t/year)..
@@ 6 Conclusion
This paper presented ICAR (Image Complexity-Aware Retrieval), the first adaptive computation method for image-text retrieval that addresses the cross-modal alignment challenge through unified embedding spaces. At the core of ICAR is the ICC module, a lightweight image complexity classifier powered by ConvNeXt-IC which predicts whether an image is simple or complex and routes simple images to early exits while complex images use full processing. During training, both the early-exit and full paths are jointly supervised via dual-path contrastive learning so that all vision embeddings remain compatible with the text embedding in a single unified space. By recognising complexity assessment as a classification task, ConvNeXt-IC achieves state-of-the-art performance (0.959 PCC and 0.956 SRCC, 4.4 faster complexity prediction), and ICAR’s dual-path training enables 20% faster image encoding with maintained retrieval quality.
Future work could enhance the approach through dynamic exit selection based on intermediate confidence scores, end-to-end integration of complexity assessment within the vision-language learning objective, and production-grade optimisations to bridge the gap between GFLOP-estimated (1.41) and measured (1.195) image encoding speedups. Additional work includes evaluating other backbones and patch sizes, and studying routing calibration under distribution shift.
6.0.1 Data and code availability
To facilitate reproducibility and future research, all code, trained model weights and data are publicly available at https://github.com/MikelWL/ICAR.